SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jan.24, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Dan Walker

Operations

  • There were no L1 accepts, but L3 went on normal during some cosmic runs on Friday swing shift (Jan.20th). It turns out to be a problem with the DCH/dataflow software.
  • There is a monthly Trigger status report after the detector operation meeting this Wednesday at IR2. A link to the report is provided here.
  • Next week's TRG operation meeting will be moved to Cedar (one time location change only).

EMT

  • No difference seen in the deadtime due to EMT for the prescaled data runs yet. Very little data with low lumi were taken for the last week after the prescale=3 factor was in. Dan is working on creating physics keys with different prescale factors. We noticed that Walt put in the FCT deadtime variables in OdcProd at the end of July last year, but they were not in the ambient archive until early October 2005.
  • Dan showed a plot for the bytes per ROM for ROMs with L1EmtSlotTC data. It looks like we are shipping 200 bytes per ROM for an empty/prescaled event. Rainer mentioned that he had asked Chris O'Grady to generate a table for each detector subsystems to see how much time the ROMs take for DAQ.
  • Dan showed a plot for the number of suppressed event (bin 0) and the number of unsuppressed event (bin 1).
  • The EMT prescale monitoring will be in the ps archive soon.

DCT/DCZ

  • Jamie and Selina tested 2 versions of Marc's new TSF firmware at the test stand. Marc has put in more diagnostics in the header to help understand the firmware better. Both versions still have missing segments and also failed MC tests at the test stand. It's possible that Marc will come to SLAC for about a week or so to work on the firmware issues.

DCZ online

  • Fang is continuing to clean up the online code. See here for an example. She tested the new and old code at the test stand yesterday, but kept getting damage 0x2 (dropped contributions) perhaps due to network problems. She pointed out that the poolsize (the amount of available memory reserved for ROM cache) is set to 8. Rainer suggested that the poolsize might need optimization if batching is active.
  • Fang updated the DCZ online lib to release 13.9.0, and some cosmic runs were taken with the new lib even though Jim already backed out the new online installation. So, it looked like the new DCZ lib worked with the previous online release. The bottom line is that no physics data were taken with this mix of releases.

L3

  • Arik is making progress in L3 tracking tagging efficiency. He will probably present something at this meeting next time.

Data Quality and Monitoring

  • Trickle injection monitoring remains to be an issue. Only about 50% of the time pilots caught the problem and took action when trikcle injection markers were lost. The DQG wouldn't be able to find problems with the trickle injection with the existing monitoring tools.
  • Rahmat showed some stripcharts.
  • According to Kris, the Run1 and Run2 reprocessing is done.
  • Kris worked on getting only one tag for L1TFMon with or without old DCT system.

Trg Config

  • Rainer updated the configuration keys for SP8 so that we have the most representative configuration key for MC to track the run condition more closely than before.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.